### v1

# Routing and wavelength assignment in quantum key distribution networks: power control heuristics for quantum-classical multiplexing

Abstract
As quantum key distribution networks grow in size and complexity, resource assignment has become increasingly important. In passive optical networks without wavelength conversion, we need to assign a full route between origin and destination with the same wavelength from a finite set. This problem is computationally intensive and the common solution in classical optical networks is using heuristics. We adapt these heuristics to hybrid quantum networks where the quantum channel can share some of the optical links with classical channels. In this quantum-classical multiplexing, nonlinear effects can become the limiting factor in the range of the network. The signal in the classical channels can be subject to Raman Scattering or Four-Wave-Mixing and produce light in the quantum channels. While these effects are not efficient, even a single photon can ruin the quantum channel. We propose heuristics for the routing and wavelength assignment problem for hybrid quantum-classical networks with power control for the classical channels. By keeping the transmitted power to its bare functional minimum, we can reduce the interference to the quantum channels. We study their efficiency under different scenarios.
Index Terms:
Routing and wavelength assignment, QKD networks, Quantum-classical multiplexing.
I Quantum key distribution networks
Communication protocols that use the laws of Quantum Mechanics allow new applications beyond classical possibilities. In particular, the limitations of quantum measurement, where the state of a superposition is destroyed after being observed, can be used to establish an arbitrarly long secret key between two parties that use a private, but insecure, quantum communication channel and a public, but authenticated, classical communication channel. This offers an additional level of security based on the physical principles that can be combined with the usual cryptographic methods based on computational complexity. Quantum Key Distribution, QKD, is a mature quantum technology for which there are multiple protocols, experimental proofs of concept, standards and commercial products [1]. Now that QKD links between two sides can be deployed with relative ease, there is a growing interest in QKD networks with multiple nodes in which any pair of nodes can establish a secret key securely.
As in most quantum communication demonstrations, in order to cover large distances fast, photons are the quantum system of choice. The main proposals are free-space QKD, including satellite communication, and optical fiber QKD networks that take advantage of the existing optical communication network infrastructure. We will concentrate on these networks.
While the first QKD demonstrations used dedicated dark fiber links, the current objective is a gradual integration with existing fiber communication networks. Ideally, QKD networks would be part of flexible Software Defined Networks, where there is a logical control plane deciding which connections can be established and how the network resources are assigned [2].
QKD links need a classical channel for public discussion, including error correction and privacy amplification, and a quantum channel which can basically use most of existing classical optical fiber technology with the important exception of amplifiers. The same principles that prevent an eavesdropper learning a quantum state forbid noiseless amplification and there is a distance limit on QKD. Photons can get lost during transmission and, above a certain loss threshold, we can no longer guarantee the observed imperfections are not due to an adversary that is attacking the system. Current QKD systems, depending on the quality of the detectors or the optical fiber used, can cover distances from tens to hundreds of kilometers. The systems that use standard optical fiber devices are limited to the smaller end of that range.
QKD hardware is chosen to minimize noise and losses. We will concentrate on discrete variable QKD protocols where the signal sent into the channel usually has less than one photon on average. This makes eavesdropping harder as we cannot measure the photon without disturbing its state and we cannot keep a copy for measurement and forward another photon. However, this means that stray photons in the quantum channel band can introduce fatal errors.
Thermal photons are not much of an issue compared to the noise photons that appear in hybrid networks where the same optical fiber carries the classical and the quantum channels needed in QKD or when we share the resources with classical traffic. In these cases, there can be contamination from the classical signal. For wavelength division multiplexing, the main contamination mechanism are nonlinear effects that convert photons from the classical channel bands into the quantum band. In particular, depending on the chosen wavelengths, Raman Scattering and Four-Wave-Mixing can create noise photons from classical channels with moderate powers with a high enough probability to be the limiting factor to the maximum link length [3, 4]. These effects are more important for longer shared paths between the classical and the quantum signals.
We propose routing heuristics that assign paths and wavelengths to the classical and quantum channels taking this contamination into account. The algorithms are inspired by the power control methods used in wireless sensor networks to reduce the consumed power and make battery lives longer [5, 6].
II Network description and traffic demands. Routing and wavelength assigment
The most natural way to represent optical networks is using graphs where each end user is at the nodes and optical fiber links are represented as the links, or edges, of the graph. We assume these links can carry a maximum total number of wavelengths $W_{T}$.
We consider a series of network demands where we are given an origin and a destination which ask for a lightpath: a path connecting the origin and destination nodes using a dedicated wavelength. Two lightpaths sharing a link must have a different wavelength, but non-overlapping lightpaths can use the same or different wavelengths. These restrictions model what happens in optical networks without wavelength conversion, which is a reasonable scenario for our QKD network.
We are interested in optimal resource assignment so that the limited number of transmitters and receivers at each node and the fibers in our network can fulfill the maximum possible number of traffic requests.
For a given topology and a series of traffic requests, the routing and wavelength assigment, RWA, problem consists in finding lighpaths for each origin and destination minimizing the number of requests that cannot be fulfilled (for instance if there are no available wavelenghts in any route between the nodes). This problem is known to be NP-complete for its usual formulation [7]. Full optimization is quite involved and, in practice, there is a set of efficient heuristics which approximate the best solution faster and using less comptuational resources [8].
A usual strategy is using the heuristic shortest-path-first-fit, SPFF, and its variations. In the heuristic, we begin using Dijkstra’s algorithm to find the shortest path between origin and destination (or the $k$ shortest paths in many cases). Then, assuming we have an ordered list of available wavelengths, we assign a lightpath to the shortest route which has a free wavelength, assigning the wavelength with the smallest index in our ordering. If no such a path can be found, the connection is blocked (the traffic request is denied).
The RWA and related problems are also relevant to QKD networks and there is an increasing interest in resource allocation methods for different kinds of quantum network [9]. We continue here the line started in [10], which also includes a comparison to other RWA proposals for QKD. The proposed heuristics focus on reducing the problems caused by multiplexing classical and quantum channels. The ultimate purpose is to give methods as simple as possible which can be deployed with little changes in current optical fiber networks.
III Channel model
When routing classical and quantum channels through the same optical fiber, we need to estimate how much noise the classical channels creates in the quantum channels. We will use a simplified model that gives an approximate measure of the expected interference without a complete, and costly, physical simulation that takes into account all the fiber parameters and has accurate models of the relevant nonlinear effects. This way, the heuristics can take into account the limiting noise effects in quantum-classical multiplexing while they are still simple and use few computational resources.
The quantum channels in a QKD network have new specific restrictions. Even if we have all the necessary resources, any quantum lightpath with excessive losses or noise will not be useful for key distribution. Above a certain Quantum Bit Error Rate, QBER, QKD protocols can no longer guarantee the generated keys are private. This limit is usually given as a function of distance, which determines the channel loss. The concrete function relating losses and the maximum achievable key rates depends on the particular protocol used, but there is always a distance above which the quantum communication phase is useless.
We will use our basic noise estimations to determine a new blocking criterion for the quantum channels. A quantum request can be blocked because there are no available wavelengths, just as classical channels are, but we will consider that quantum channels that cannot generate secret key bits above a certain rate must be blocked to avoid them hogging the network resources without a useful outcome.
We will use the quantum signal-to-noise ratio, QSNR, in the quantum channel as a guide. In most natural scenarios, a smaller signal-to-noise ratio in the quantum channel is related to a higher QBER and a lower established key rate. We will use a QSNR threshold to determine whether a quantum channel can provide fresh secret bits to a key or not. For each QKD protocol, we can always find out which SNR threshold corresponds to an exchange that is doomed to fail due to a QBER above the minimum value. In general, this is a direct function of the total distance of the quantum link, but, in hybrid QKD networks, the classical part and the noise it induces is also relevant.
III-A Classical and quantum signal to noise ratios
We will start with general definitions of the signal-to-noise ratios we need that will be later refined in the model.
In the classical channels we use the usual definition for the signal-to-noise ratio
$SNR=\frac{P}{N}$
as the ratio between the signal power $P$ and the noise power $N$. For values below a certain threshold, which depends on the receivers, we assume the signal cannot be interpreted correctly. If we have a fixed power at each transmitter, the receivers at the end of shorter links will receive a higher power. We are going to assume there is a fixed noise power for each link, mostly related to thermal noise at the receivers.
For the usual lengths in QKD, most classical channels can be transmitted at reasonable power levels. This means that, for short paths, we are actually using more power than required for correct classical communication. Apart from increasing the energy bill, this excess transmission power harms the quantum channels.
The heuristics we propose for routing will adjust the classical power at the transmitter so that there is a fixed power for any link at the receiver after a route has been chosen. The power can be adjusted to whatever $SNR$ threshold is considered acceptable. This reduces the total transmitted power which will, in turn, reduce the number of noise photons in the classical channel.
We estimate the effects on the quantum channel by defining a quantum signal-to-noise ratio
$QSNR=\frac{P_{q}}{N}=\frac{P_{q}}{N_{f}+N_{c}}=\frac{P_{o}10^{-\alpha L}}{N_{f}+\sum\gamma_{nl}P_{s}}.$
We consider all the powers are normalized to the single photon level for the corresponding wavelength so that we can interpret them as a photon number. $P_{q}$ gives an estimation of the expected photon number that reaches the receiver in a detection window. It can be computed from the transmitted photon number and the attenuation in the journey to the destination node. We will also include in that term effects that are equivalent to losses, like detection efficiencies below 1, which are common for the simplest single photon detectors. For instance, many systems use avalanche photodiodes which are notoriously inefficient at the IR telecom bands, where 10-20% efficiencies are the norm.
The normalized noise power can be interpreted as the average number of spurious photons. We include two terms for general noise, $N_{f}$, and noise caused by the classical channels $N_{c}$.
The fixed amount $N_{f}$ that collects thermal photons and equivalent effects, like dark counts in the detectors (false detections caused by phonons or other effects that are not related to the incoming light).
We will give a simplified model to estimate $N_{c}$ as the product of a non-linear coefficient $\gamma_{nl}$ by a sum of the power of the classical power $P_{s}$ of each classical channel that shares a link with the quantum channel normalized by the shared total distance. This is a linearization that does not cover the full physical interaction, but it can give a reasonable estimation of which classical channels are likely to introduce the most noise without a detailed physical model of each situation. The parameters can be adjusted to model each particular scenario, including using different non-linear coefficients for different links using different optical fibers or whether the classical and quantum signals are copropagating or counterpropagating.
IV System Model
We can give a formal description of the QKD networks we want to study under a few reasonable assumptions. First we will describe the exact problem we want to solve and then we will choose a relevant scenario where we perform our numerical experiments.
IV-A Problem Statement
We consider an arbitrary physical topology of an optical network as a directed graph $G(N,L,W_{T},W_{Q})$ composed of $N$ nodes and $L$ optical fiber links. $W_{T}$ is the total number of wavelengths in each link, while $W_{Q}$ is the number of reserved wavelengths for the quantum channel. We consider quantum requests $R(s,d)$ where $s$ is the source node and $d$ is the destination node. In the experiments, for each quantum request we require that four channels are established: the quantum channel, two classical control channels (one from $s$ to $d$ and one from $d$ to $s$), and an additional classical channel for user data transmission.
We want to provide a valid lightpath for all the traffic requests for the given topology. In our hybrid networks, the signals in the classical channels can disrupt the quantum transmission. We will propose RWA heuristics using power control over the classical channels to minimize these interferences from the classical into the quantum lightpaths and compare them to the usual classical heuristics.
IV-B Network Assumptions
We consider optical networks with neither wavelength converters nor amplifiers. To serve a connection request between any pair of nodes, not necessarily adjacent in the network, we establish a lightpath to which a route and a wavelength are allocated. Lightpaths whose routes share at least one link cannot be allocated the same wavelength.
We assume a QKD software-defined network (QKD SDN) with a classical control plane which uses the proposed algorithms to evaluate and allocate network resources to all the connection requests.
For our experiments, we generate random network topologies with a minimum of $8$ nodes and a maximum of $12$ nodes. The topologies are created so that any node has at least one neighbour to avoid isolated nodes in the graph. Two nodes are physically connected with a given link probability. When two adjacent nodes are connected, we set two unidirectional links (one from $a$ to $b$ and one from $b$ to $a$). We have set relatively short links of tens of kilometers due to restrictions in the achievable secure distance for commercial QKD links.
We assume static traffic, i.e., the connection requests are known in advance, although the control plane evaluates each request sequentially. The RWA algorithm considers the state of the network and uses the available resources to establish lightpaths. We consider two types of requests:
  • QKD channel demands. This request requires establishing a quantum channel between a source and a destination node to generate a common secret key. The request will be associated to four lightpaths: a quantum channel to send the quantum states, two classical channels for bidirectional communication between the pair of nodes for synchronization and authenticated public discussions, and a classical channel to send the encrypted information with the generated key. Each channel requires a dedicated lightpath.
  • Classical channel demands. We also consider the user can ask for purely classical connections on top of the QKD network.
The performance of the algorithms is measured in terms of blocking probability, which is the ratio between the rejected connections and total demands. Connections for both quantum and classical requests are rejected if the algorithm cannot find a path and a wavelength between the source and destination nodes. Additionally, QKD lightpaths should be established only if they allow the generation of a secret key that meets a certain quality parameter, which we measure in terms of Quantum Signal-to-Noise Ratio (QSNR), which will be introduced in the following subsection. Furthermore, if any classical channel, either associated with the requested quantum channel or with a purely classical demand causes noise to any established quantum channel that introduces more errors than the protocol can tolerate, that connection should not be established, since it will cause a failure in the key generation procedure while wasting network resources.
Similarly, if the noise from classical existing signals would introduce into a quantum channel more errors than the protocol can tolerate, even if we have available wavelengths, the connection should not be established. This is because the key generation procedure will inevitably fail, leading to wasted network resources. To address this, we define a conservative threshold for the QSNR for each quantum lightpath, so that the connection is blocked if its QSNR falls below this threshold.
IV-C Transmission Power Control and RWA Algorithms in QKD Networks
IV-C1 Transmission Power Control
We adapt RWA algorithms described in [10] by adding an adaptive power launch feature for the classical channels. The RWA algorithms find a candidate path and available wavelength and calculate the power launch for that particular path, so that the signal satisfies a given signal-to-noise ratio (SNR) at the receiver, following a similar procedure to [11, 12]. The SNR equation can be expressed as:
$\text{SNR}=\frac{S_{c}}{N}=\frac{10^{-\alpha_{c}L_{c}}P_{tx,c}}{N}$ (1)
where $\alpha_{c}$ is the attenuation at the transmission wavelength, $L_{c}$ is the total length of the candidate path, $P_{tx,c}$ is the launch power of the classical signal and $N$ is the noise term. All the signal and noise powers will be normalized to a reference level. From this expression, we can calculate the launch power as:
$P_{tx,\ c}=\frac{SNR\cdot N}{10^{-\alpha_{c}L_{c}}}.$ (2)
With these calculations, the control plane can adjust the power of the classical signal at each source transmitter to a level that both guarantees a clear reception and minimizes interference into the quantum channels.
IV-C2 RWA Algorithms
We study three different heuristics, both with and without power control.
  • Minimum Quantum Distance Overlap (MQDO): This algorithm identifies all possible paths between a pair of nodes, marking the links of each path that are shared with established quantum channels. Then, the algorithm attempts to allocate an available wavelength on the path with the lower accumulated shared distance using the first-fit technique. If the allocation fails, the algorithm proceeds with the next path in ascending order of accumulated shared distance with established quantum channels and tries again. The process continues until finding a suitable wavelength or until all paths are exhausted. If the allocation is not possible, the connection is blocked.
  • Minimum Quantum Classical Channel Overlap (MQCCO): Similar to MQDO, the algorithm calculates the accumulated shared distance with quantum connections for each possible path, considering other established classical channels sharing any link of the path. Specifically, for each established quantum channel, if a link of the candidate path is marked as shared, and that link is also traversed by an established classical channel, the contribution to the accumulated shared distance is doubled. After that, the algorithm proceeds as described in MQDO. It tries to allocate an available wavelength to a candidate path using the first-fit technique. If the allocation is not possible, the algorithm moves to the next path in increasing order of shared distance. If the algorithm is unable to allocate resources after evaluating all candidate paths, the request is blocked.
  • Quantum Totally Disjoint (QTD): The algorithm attempts to allocate a path and an available wavelength that do not share links with any established quantum channel. If the algorithm cannot find such a path or available spectral resources, the connection is blocked.
All these proposals are compared to the reference classical RWA algorithm $k$-Shortest Paths (KSP) and First Fit (FF) [8] for both the quantum and the classical channels, integrating the transmission power control protocol for the classical channels. The algorithm computes the $k$ shortest paths between the source and destination. Subsequently, in increasing order of path length, the algorithm attempts to allocate the first available wavelength along each path until a suitable pair of path-wavelength is found. If the algorithm cannot allocate the required resources, the request is blocked.
For the quantum channel, all the proposed algorithms verify if the selected path meets the requirements of QSNR. In the case of quantum requests, if the resulting QSNR falls below a certain threshold, the algorithms sequentially evaluate the next path until they identify a path with a free wavelength that satisfies the QSNR threshold or block the request otherwise.
Conversely, the establishment of classical channels must not degrade the performance of existing quantum connections. Hence, the RWA algorithms ensure that the selected path does not take the QSNR of established quantum channels below the specified QSNR threshold. If a path fails to meet this criterion, the algorithms proceed to assess subsequent paths until finding one that does not degrade the established quantum channels. If no such path is found, the connection is blocked. The QSNR is estimated with:
$\text{QSNR}\ =\frac{10^{-\alpha_{q}L_{q}}P_{tx,q}}{N_{F}+\gamma_{nl}\sum_{i}P_{tx,i}\cdot e^{-\alpha_{c}L_{i}}}$ (3)
where $\alpha_{q}$ is the attenuation of the fiber at the transmission wavelength of the quantum signal, $L_{q}$ is the length of the route allocated to quantum channel $q$, $P_{tx,q}$ is the transmitted power of the quantum channel, $N_{F}$ is a fixed noise term, $\gamma_{nl}$ is a fixed nonlinear coefficient term that can be adapted to the physical parameters of the fiber used in the studied network, $P_{tx,i}$ represents the classical signal launch power at the beginning of the shared span $i$, with shared length $L_{i}$ and $\alpha_{c}$ represents the attenuation coefficient at the transmission wavelength of the classical channel. We use two metrics to calculate the shared span:
  1. 1.
    Actual shared length: $L_{i}$ represents the actual length of the shared fiber.
  2. 2.
    Effective length: if $L$ represents the actual distance in span $i$ for which the power can be assumed as constant, $L_{i}$ will be set to the effective distance which is computed as [13]:
    $L_{i}=L_{\text{eff}}=\frac{1-e^{-\alpha_{c}L}}{\alpha_{c}},$ (4)
for an $\alpha_{c}$ given in natural units. This distance takes into account that the attenuation of the classical signal during propagation implies a reduced number of noise photons created by nonlinear effects. $L_{\text{eff}}$ is the distance which produces the equivalent nonlinear effects for a fixed signal power.
V Simulation Set up and Results
V-A Set up
We investigate the effects of the transmission power protocols on $10$ randomly generated topologies. The number of nodes for each topology is randomly chosen between $5$ and $10$. Each pair of nodes is connected via two unidirectional fibers with a probability of $0.5$. To ensure network connectivity, each node must have at least two neighbours. The length of the fiber links are randomly chosen between $10$ and $20$ km. Examples of the generated topologies are shown in figures 1,2.
Figure 1: Example of a random topology with $6$ nodes generated in this study.
Figure 2: Example of a random topology with $9$ nodes generated in this study.
[…]
We also normalize the launch power of the classical channels with respect to the maximum launch power. This maximum power launch is determined by ensuring that the longest possible path for that particular connection achieves a minimum SNR of $20$ dB for a normalized noise power $N=1$. Once the RWA algorithms identify the candidate path, they compute the required transmission power for that particular path and normalize it to the maximum launch power. Consequently, both $P_{tx,q}$ and $P_{tx,c}$ are normalized at their respective scales. Since the transmission power of quantum and classical lightpaths differ in magnitude, this difference is accounted for in the $N_{c}$ term.
We use the request blocking ratio and the average SNR as the figures of merit to evaluate the performance of the algorithms. The total number of requests arriving to the network varies from $10$ to $100$ with an increment of $10$ requests. For each topology, $500$ simulations are run. Figures include $95\%$ confidence intervals. The averages in the experiments are fairly stable and the confidence intervals are difficult to appreciate in some of the figures.
V-B Results
Figure 3: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the actual shared fiber length $L_{i}$.
Figure 4: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the shared effective length $L_{\text{eff}}$.
Figure 5: Difference in the SNR achieved by the RWA algorithms when the Power Control protocol is applied, considering the shared effective length.
Figures 3 and 4 show the blocking ratio of the algorithms with transmission power control compared to the same RWA algorithms that do not apply this control protocol, using $L_{i}$ and $L_{\text{eff}}$ respectively. It can be observed that the results are very similar for both shared distance calculations. Therefore, we will analyze the results obtained using $L_{\text{eff}}$.
Among the algorithms with no Transmission Power Control protocol, KSP-FF demonstrates the best performance in terms of accepted requests, achieving a $0\%$ rejection rate for total requests under $50$ and outperforming the quantum-oriented MQDO and MQCCO by up to $2$ orders of magnitude. QTD exhibits the highest blocking ratio, with blocking ratios above $10^{-1}$, as it consumes the most resources. When the Power Transmission Protocol is applied, all the algorithms except QTD show improved performances. This is expected. In QTD no fiber link shares quantum and classical channels. KSP-FF reduces the blocking ratio by up to $2$ orders of magnitude, while MQDO and MQCCO can improve their performance by up to $1$ order of magnitude. Therefore, effective control of the power transmission of the classical channels enhances the performance of the QKD network.
Figure 5 shows the achieved QSNR of the established quantum channels. The QSNR is consistently above the $15$ dB threshold across all algorithms. Without active transmission power control, the QSNR decreases, as expected, due to increased noise on the quantum lightpaths caused by the higher transmission power of the classical channels. A smaller QSNR is related to a higher QBER which, in turn, means the QKD protocol has a slower key generation rate. Noisier channels produce secret key bits more slowly.
Specifically, KSP-FF achieves a QSNR between $24-27$ dB without power control and improves by up to $3$ dB for the higher number of requests when power control is active. MQDO and MQCCO exhibit similar behavior, with improvements up to $2$ dB. QTD behaves the same with and without power control (there are no shared paths).
When power control is not active, KSP-FF yields lower quality channels compared to quantum-oriented RWA algorithms. While quantum-oriented algorithms can enhance the quality of quantum channels by up to $2$ dB, they do so at the cost of rejecting more requests than KSP-FF. However, with active power control, the quality of channels established by the algorithms is nearly identical.
These experiments suggest proper control of classical transmission power can enhance signal quality and enables the use of classical RWA algorithms for solving the resource allocation problem.
V-C Performance of mixed traffic scenarios.
We also investigate the power control protocol over RWA algorithms in mixed traffic scenarios, i.e., networks in which both quantum and classical requests are present. We consider a network trying to establish QKD links over a network with classical traffic outside our control. Specifically, we analyze the performance of the network for $90$ total requests. We define the classical traffic load as the fraction of the total requests that belong to purely classical traffic. This classical load varies from 00 (representing all the request are quantum) to $1$ (representing all the requests are classical). All other simulation parameters remain unchanged and results are represented with $95\%$ confidence intervals.
Figure 6 shows the blocking ratio of the RWA algorithms with power control protocol and when this protocol is not active. As in 4, when the classical traffic load is 00, the best performing algorithm is KSP-FF with power control protocol. However, when the classical traffic load increases, the performance of KSP-FF with and without power control deteriorates. This decline is due to the potential interference introduced by the increasing number of classical lightpaths, which degrades the QSNR of the established quantum channels. The algorithms are designed to cause a rejection of the classical connection request if they degrade the quality of the established quantum channels below a certain threshold, hence the increase on the blocking ratio.
On the other hand, MQDO and MQCCO improve their performance because they are designed to minimize the shared distance with established quantum channels, thereby reducing interference and degradation of these lightpahts’ quality. Considering that each quantum connection requires the establishment of three additional classical lightpaths, an increase in the classical load implies a reduction in the total number of quantum requests and the associated classical connections, leading to a decrease in the network resource consumption and, consequently, a reduction of the blocking ratio compared to KSP-FF by almost $2$ orders of magnitude. Moreover, applying the power control protocol, we can reduce the blocking ratio by an order of magnitude compared to the same algorithms without power control. Finally, the worst-performing algorithm is QTD, although for the higher number of classical requests, finding a disjoint path for the classical channels can improve the blocking ratio with respect to KSP-FF.
The QSNR achieved by these algorithms is shown in figure 7. The power control protocol enhances the performance of KSP-FF, MQDO, and MQCCO by up to $2$ dB, and QTD by almost $1$ dB (fot power control restricted to the classical channels in the quantum requests). In QKD networks with classical traffic, KSP-FF with power control can slightly improve the quality of the established channels, although at the expense of increasing the blocking ratio. Also, MQCCO and MQDO achieve almost identical QSNRs for the established quantum channels and equal blocking ratio, meaning that the extra effort in accounting for all the established classical channels that MQCCO makes when selecting candidate paths does not provide benefits in terms of either blocking ratio or quality of the quantum channel.
Figure 6: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the actual shared fiber length $L_{i}$.
Figure 7: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the shared effective length $L_{\text{eff}}$.
VI Conclusions
In this study, we evaluate the performance of a power control algorithm over a traditional RWA algorithm like KSP-FF and other quantum-oriented RWA algorithms proposed in the literature. This protocol aims at reducing the launch power of the classical lightpaths in order to reduce the noise introduced to the quantum connections. Our study accounts for the undesired noise that classical channels potentially introduce into the quantum channels, proposing a QSNR formula inspired by the SiNR in wireless networks. The study demonstrates that implementing a power control protocol effectively mitigates the noise introduced to quantum connections. This approach enhances the QSNR by reducing interference, thereby improving the quality of quantum channels.
We have found that, in QKD networks, the application of the power control protocol alone can significantly improve the blocking ratio and enhances the quality of the established channels. This finding underscores the protocols’s efficiency in quantum communication networks, where minimizing the interference of the classical channels is critical for maintaining high-quality secure communication.
Lastly, in mixed traffic scenarios, where both quantum and classical connection requests are present, the study finds that the increasing number of established classical lightpaths can degrade the quality of the quantum channels despite the use of power control. Quantum-oriented RWA algorithms, which consider the unique requirements of quantum communication, are necessary to effectively manage network resources and maintain high QSNR. These algorithms are crucial for minimizing the impact of classical channels on quantum communication, ensuring the reliability and security of the network.
VII Acknowledgments
Lidia Ruiz has been funded by the European Union NextGeneration UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León. J.C. Garcia-Escartin has been funded by the European Union NextGeneration UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León and the Spanish Ministerio de Ciencia e Innovación (MCIN), project PID2020-119418GB-I00.